# The Surprising Effectiveness of Multimodal Large Language Models for Video Moment Retrieval

### v1

@@ Abstract
Recent studies have shown promising results in utilizing multimodal large language models (MLLMs) for computer vision tasks such as object detection and semantic segmentation. However, many challenging video tasks remain under-explored. Video-language tasks necessitate spatial and temporal comprehension and require significant compute. Therefore, prior works have developed complex, highly specialized architectures or leveraged additional input signals such as video transcripts to best encode contextual and temporal information, which limits their generality and can be impractical. One particularly challenging task is video moment retrieval, which requires precise temporal and contextual grounding. This work demonstrates the surprising effectiveness of leveraging image-text pretrained MLLMs for moment retrieval. We introduce Mr. BLIP[Uncaptioned image] (Mr. as in Moment Retrieval), a multimodal, single-stage model that requires no expensive video-language pretraining, no additional input signal (e.g., no transcript or audio) and has a simpler and more versatile design than prior state-of-the-art methods. We achieve a new state-of-the-art in moment retrieval on the widely used benchmarks Charades-STA, QVHighlights, and ActivityNet Captions and illustrate our method’s versatility with a new state-of-the-art in temporal action localization on ActivityNet. Notably, we attain over 9% (absolute) higher Recall (at 0.5 and 0.7 IoU) on the challenging long-video multi-moment QVHighlights benchmark. Our code is publicly available  [^1] https://github.com/sudo-Boris/mr-Blip
@@ 1 Introduction
In this work, we develop Mr. BLIP[Uncaptioned image], an image-text pretrained, multimodal large language model based on BLIP-2 [37] and a novel multi-modal input sequence. The model design is capable of interpreting video data, and leverages interleaved temporal grounding of the input video signal (in the form of time-stamps) to enhance its predictions. As illustrated in Fig. 1, we also start of with both visual frames and query. Prior work (bottom), often leverages video features [20] and CLIP [51] to then train a complex, task-specific feature fusion module that predicts a fixed set of candidates. In contrast, Mr. BLIP (top) leverages a sampled subset of the frames directly and utilizes the MLLM’s attention mechanism to learn the contextual and temporal relationship between visual frames, their interleaved temporal grounding, and the respective natural language query. Later in the paper, we discuss the relevance of deliberate design choices by conducting extensive ablation studies and demonstrate the surprising effectiveness of leveraging image-text pretrained MLLM for moment retrieval. Our approach achieves state-of-the-art results across the widely used benchmarks, Charades-STA [21], QVHighlights [34], and ActivityNet Captions [30]. Moreover, to illustrate our model’s versatility, we evaluate our method on a second task, Temporal Action Localization (TAL) on ActivityNet 1.3 [4], and achieve a new SoTA.
In summary, our main contributions are the following: (i) We introduce Mr. BLIP for moment retrieval, leveraging image-text pretrained MLLMs by casting the moment retrieval task as an open-ended sequence-to-sequence problem. (ii) To enhance the temporal understanding of events in input videos, we design a novel multimodal input sequence. (iii) Our Mr. BLIP model improves the state-of-the-art across the most widely used moment retrieval benchmarks [21, 34, 30] and temporal action localization [4]. (iv) Finally, extensive experiments and ablations demonstrate the effectiveness of Mr. BLIP and its design choices. We find that representing time as absolute seconds and interleaving it with the respective frames yields the best performance.
@@ 2.1 Moment Retrieval Models
Proposal-free works include regression-based methods [34, 60, 41, 48, 65] that involve learning the interactions between both modalities to directly predict the temporal boundaries of a relevant moment without the need for a first stage to generate initial proposals. Nevertheless, proposal-free methods still predict a fixed number of candidate windows alongside a confidence score. The predicted windows are sorted based on the confidence score and are evaluated with proxy metrics based on different overlapping thresholds such as Recall@{0.5, 0.7}.
@@ 4.1 Experimental Setup
We validate Mr. BLIP on the three most widely used video moment retrieval (MR) datasets and illustrate our model’s versatility on the most common temporal action localization (TAL) benchmark:
@@
ActivityNet Captions [30] contains 20,000 videos with an average duration of 2 minutes. The dataset contains 100,000 segments where each was human-annotated with a caption that contains on average 13.5 words. The dataset is divided into three splits, train (10,024), val_1 (4,926), and val_2 (5,044). Following [60], we use the train split for training, val_1 for validation, and val_2 for testing.
ActivityNet 1.3 [4] uses the same 20,000 untrimmed videos and the same split ratio as [30] but has 200 action class labels associated with a specific moment in the video as annotations. Specifically, most videos contain only one labeled segment but can have multiple; if there are more than one, each segment is from the same action class. We follow standard practice [60, 40, 8], train our model on the training set, and report the results on the validation set.
@@
The most commonly used metrics for MR are Recall@K and mean average precision (mAP) computed under different Intersection over Union (IoU) thresholds. The Recall@K metric is defined as the percentage of the top-K predicted segments having a larger temporal IoU than the threshold with a ground truth segment. Following the recent development [33, 34, 46, 47], we report the more challenging $R1@0.5$ and $R1@0.7$ scores, which correspond to the $Recall@1$ scores at IoU thresholds of $0.5$ and $0.7$, respectively, and the mean IoU (mIoU) score. Following prior work, for MR on QVHighlights and TAL on ActivityNet 1.3, we report the mAP score at IoU thresholds of $0.5$ and $0.75$, respectively.
@@
For Charades-STA, we extract 20 video frames for each video, which is significantly less than prior work which requires 128 frames [60]. We train for up to 50 epochs with a batch size of 32 samples split across 4 A100-80GB GPUs for about 55 GPU hours. For QVHighlights and ActivityNet Captions and 1.3, we extract 60 video frames for each video, again less than required by prior work. For TAL on ActivityNet 1.3, we augment the class names using Kinetics prompts released by [51], e.g. “a video of a person performing {label}”. We train for up to 50 epochs with an effective batch size of 32 samples by splitting them across 8 A100-80-GB GPUs and leveraging accumulated gradients set to 4 (i.e., we have one sample per GPU), and train for about 170 GPU hours.
@@ 4.2 Ablation Studies
By default, our Mr. BLIP models for Charades-STA (QVHighlights/ ANet Captions/ ANet 1.3) leverage 20 (60) visual frames, their associated timestamps, video duration, a query, and a task prompt. If not stated otherwise, we represent time as simple text tokens of whole seconds and interleave these with frame tokens. In the following, we ablate the impact of these design choices on the downstream moment retrieval performance by reporting results on our Charades-STA validation set.
@@ 4.3 Comparison to the State-of-the-Art
Table 4: Comparison with state-of-the-art. Charades-STA [21] and QVHighlights [34]. CLIP: UnLoc [60] pretrains the backbone with action classification datasets Kinetics 400/700. SF: Slow-Fast backbone. Note, “-” means the result is not known, while grey means the result can not be obtained (no standard validation set available). Discussion in the main text.
Charades-STA QVHighlights
Method Backbone R1@.5 R1@.7 mIoU R1@.5 R1@.7 mAP@.5 mAP@.75
Validation set ([25, 60] only provide results on the validation set for QVHighlights)
@@
UniVTG [39] SF+CLIP 58.01 35.65 50.10 58.86 40.86 57.60 35.59
QD-DETR [47] SF+CLIP 57.31 32.55 62.40 44.98 62.52 39.88
CQ-DETR [46] SF+CLIP 58.44 36.34 50.13 65.43 48.38 64.51 42.77
@@
Mr. BLIP[Uncaptioned image] BLIP-2 69.31 49.29 58.63 74.77 60.51 68.12 53.38
@@ 4.3.1 Moment Retrieval
Notably, we also outperform the best reported approach on QV-Highlights validation set, UnLoc-L[60], with far fewer input frames, 128 vs. 20, and without expensive video pretraining of the backbone.
@@
Furthermore, as shown in Table 5(a), Mr. BLIP outperforms the previous SoTA [60] in moment retrieval on ActivityNet Captions by 5.62% and 5.35% on R1@0.5 and R1@0.7, respectively.
@@
Table 5: Comparison with state-of-the-art on ActivityNet.
(a) ANet Single Sentence Moment Retrieval on ActivityNet Captions [30]
Method R1@0.5 R1@0.7
VLG [52] 46.3 29.8
UnLoc-L [60] 48.30 30.20
Mr. BLIP[Uncaptioned image] 53.92 35.55
(b) ANet Temporal Action Localization on ActivityNet 1.3 [4]
Method mAP@0.5 mAP@0.75
UnLoc-L [60] 59.30
AdaTAD [40] 61.72 43.35
Mr. BLIP[Uncaptioned image] 67.53 53.13
4.3.2 Temporal Action Localization
Finally, although our main goal is to explore leveraging an MLLM for the more sophisticated video-language task of moment retrieval, we demonstrate the versatility of our model by applying it to Temporal Action Localization on ActivityNet 1.3 as shown in Table 5(b). We again achieve a new state-of-the-art beating AdaTAD [40] by 5.81% and 9.78% on mAP@0.5 and mAP@0.75, respectively.

### v2

@@ Abstract
Recent studies have shown promising results in utilizing multimodal large language models (MLLMs) for computer vision tasks such as object detection and semantic segmentation. However, many challenging video tasks remain under-explored. Video-language tasks necessitate spatial and temporal comprehension and require significant compute. Therefore, prior works have developed complex, highly specialized architectures or leveraged additional input signals such as video transcripts to best encode contextual and temporal information, which limits their generality and can be impractical. One particularly challenging task is video moment retrieval, which requires precise temporal and contextual grounding. This work demonstrates the surprising effectiveness of leveraging image-text pretrained MLLMs for moment retrieval. We introduce Mr. BLIP[Uncaptioned image] (Mr. as in Moment Retrieval), a multimodal, single-stage model that requires no expensive video-language pretraining, no additional input signal (e.g., no transcript or audio) and has a simpler and more versatile design than prior state-of-the-art methods. We achieve a new state-of-the-art in moment retrieval on the widely used benchmarks Charades-STA, QVHighlights, and ActivityNet Captions. Notably, we attain over 9% (absolute) higher Recall (at 0.5 and 0.7 IoU) on the challenging long-video multi-moment QVHighlights benchmark. Our code is publicly available  [^1] https://github.com/sudo-Boris/mr-Blip.
@@ 1 Introduction
In this work, we develop Mr. BLIP[Uncaptioned image], an image-text pretrained, multimodal large language model based on BLIP-2 [36] and a novel multi-modal input sequence. The model design is capable of interpreting video data, and leverages interleaved temporal grounding of the input video signal (in the form of time-stamps) to enhance its predictions. As illustrated in Fig. 1, we also start of with both visual frames and query. Prior work (bottom), often leverages video features [19] and CLIP [49] to then train a complex, task-specific feature fusion module that predicts a fixed set of candidates. In contrast, Mr. BLIP (top) leverages a sampled subset of the frames directly and utilizes the MLLM’s attention mechanism to learn the contextual and temporal relationship between visual frames, their interleaved temporal grounding, and the respective natural language query. Later in the paper, we discuss the relevance of deliberate design choices by conducting extensive ablation studies and demonstrate the surprising effectiveness of leveraging image-text pretrained MLLM for moment retrieval. Our approach achieves state-of-the-art results across the widely used benchmarks, Charades-STA [20], QVHighlights [33], and ActivityNet Captions [29].
In summary, our main contributions are the following: (i) We introduce Mr. BLIP for moment retrieval, leveraging image-text pretrained MLLMs by casting the moment retrieval task as an open-ended sequence-to-sequence problem. (ii) To enhance the temporal understanding of events in input videos, we design a novel multimodal input sequence. (iii) Our Mr. BLIP model improves the state-of-the-art across the most widely used moment retrieval benchmarks [20, 33, 29]. (iv) Finally, extensive experiments and ablations demonstrate the effectiveness of Mr. BLIP and its design choices. We find that representing time as absolute seconds and interleaving it with the respective frames yields the best performance.
@@ 2.1 Moment Retrieval Models
Proposal-free works include regression-based methods [33, 58, 39, 46, 63] that involve learning the interactions between both modalities to directly predict the temporal boundaries of a relevant moment without the need for a first stage to generate initial proposals. Nevertheless, proposal-free methods still predict a fixed number of candidate windows alongside a confidence score.
The predicted windows are sorted based on the confidence score and are evaluated with proxy metrics based on different overlapping thresholds such as Recall@{0.5, 0.7}.
@@ 4.1 Experimental Setup
We validate Mr. BLIP on the three most widely used video moment retrieval (MR) datasets:
@@
ActivityNet Captions [29] contains 20,000 videos with an average duration of 2 minutes. The dataset contains 72,000 segments where each was human-annotated with a caption that contains on average 13.5 words. The dataset is divided into three splits, train (37,421), val_1 (17,505), and val_2 (17,031). Following [58], we use the train split for training, val_1 for validation, and val_2 for testing.
@@
The most commonly used metrics for MR are Recall@K and mean average precision (mAP) computed under different Intersection over Union (IoU) thresholds. The Recall@K metric is defined as the percentage of the top-K predicted segments having a larger temporal IoU than the threshold with a ground truth segment. Following the recent development [32, 33, 44, 45], we report the more challenging $R1@0.5$ and $R1@0.7$ scores, which correspond to the $Recall@1$ scores at IoU thresholds of $0.5$ and $0.7$, respectively, and the mean IoU (mIoU) score. Following prior work, for MR on QVHighlights, we report the mAP score at IoU thresholds of $0.5$ and $0.75$, respectively.
@@
For Charades-STA, we extract 20 video frames for each video, which is significantly less than prior work which requires 128 frames [58]. We train for up to 50 epochs with a batch size of 32 samples split across 4 A100-80GB GPUs for about 55 GPU hours. For QVHighlights and ActivityNet Captions, we extract 60 video frames for each video, which is less than required by prior work. We train for up to 50 epochs with an effective batch size of 32 samples by splitting them across 8 A100-80-GB GPUs and leveraging accumulated gradients set to 4 (i.e., we have one sample per GPU), and train for about 170 GPU hours.
@@ 4.2 Ablation Studies
By default, our Mr. BLIP models for Charades-STA (QVHighlights/ ANet Captions) leverage 20 (60) visual frames, their associated timestamps, video duration, a query, and a task prompt. If not stated otherwise, we represent time as simple text tokens of whole seconds and interleave these with frame tokens. In the following, we ablate the impact of these design choices on the downstream moment retrieval performance by reporting results on our Charades-STA validation set.
@@ 4.3 Comparison to the State-of-the-Art
Table 4: Comparison with state-of-the-art. Charades-STA [20], QVHighlights [33], and ActivityNet Captions [29]. CLIP: UnLoc [58] pretrains the backbone with action classification datasets Kinetics 400/700. SF: Slow-Fast backbone. Discussion in the main text.
Method Backbone mIoU R1@.5 R1@.7 mAP@.5 mAP@.75
QVHighlights [33] Validation set
@@
UniVTG [38] SF+CLIP 58.86 40.86 57.60 35.59
QD-DETR [45] SF+CLIP 62.40 44.98 62.52 39.88
CG-DETR [44] SF+CLIP 65.43 48.38 64.51 42.77
Mr. BLIP[Uncaptioned image] BLIP-2 74.77 60.51 68.12 53.38
Charades-STA [20] Test set
UniVTG [38] SF+CLIP 50.10 58.01 35.65
QD-DETR [45] SF+CLIP 57.31 32.55
CG-DETR [44] SF+CLIP 50.13 58.44 36.34
@@
Mr. BLIP[Uncaptioned image] BLIP-2 58.63 69.31 49.29
ActivityNet Captions [29] Single sentence MR. Test set
VLG [50] GCNeXt 46.30 29.80
UnLoc-L [58] CLIP 48.30 30.20
Mr. BLIP[Uncaptioned image] BLIP-2 53.92 35.55
@@ 4.3.1 Moment Retrieval
Notably, we also outperform the best reported approach on QVHighlights validation set, UnLoc-L[58], with far fewer input frames, 128 vs. 20, and without expensive video pretraining of the backbone.
@@
Furthermore, as shown in Table 4, Mr. BLIP outperforms the previous SoTA [58] in moment retrieval on ActivityNet Captions by 5.62% and 5.35% on R1@0.5 and R1@0.7, respectively.